OpenStudy (anonymous):

\[\sqrt{48/75} * \sqrt{2}\]

mathslover (mathslover):

ok so do you know that : \[\sqrt{a}*\sqrt{b}=\sqrt{ab}\]

OpenStudy (anonymous):

\[\sqrt{\frac {2*2*2*2*3}{5*5*3}}*\sqrt{2}\] \[\frac {4}{5} \sqrt{\frac {3}{3}}*\sqrt{2}\] \[\frac {4}{5}*1*\sqrt{2}\] \[\frac {4}{5}\sqrt{2}\]

mathslover (mathslover):

It was just a hint : can you solve further ? May i know how u simplified it

mathslover (mathslover):

@shubham.bagrecha please let the asker do it on herself ..

OpenStudy (anonymous):

1.4142...

OpenStudy (anonymous):

ok

mathslover (mathslover):

no problem shubham .. it was just a request to u .. no problem

OpenStudy (anonymous):

ok i will remember.

OpenStudy (anonymous):

48 = 2 ^ 4 * 3. 75 = 3 * 5^2. Use the property that mathslover mentioned above to combine that under one sqrt. also, sqrt(a/b) = sqrt(a) / sqrt(b). and sqrt(a^(2*b)) = sqrt( (a^b)^2 ) = a^b.

OpenStudy (anonymous):

use those properties to solve the thing.

OpenStudy (anonymous):

k..